OpenStudy (anonymous):

solve using substitution, show work -7x - 2y = -13 x - 2y = 11

OpenStudy (anonymous):

i have. -7x - 2y = -13 -2y = -13 + 7x what do i do from here?

OpenStudy (anonymous):

how did you get -2y = 11 - x?

OpenStudy (campbell_st):

ok... so all you have to do is substitute it into the 2nd equation x + (-13 + 7x) = 11

OpenStudy (anonymous):

nothing with -2y ?

OpenStudy (campbell_st):

I just used the 2nd equation while you used the 1st equation so if I used -2y = 11 - x then substituting it into the 1st equation you get -7x + (11 - x) = -13

OpenStudy (anonymous):

i see.

OpenStudy (campbell_st):

both methods will give the same answer...

OpenStudy (anonymous):

so. x - 2 (-13 + 7x) = 11 x + 26 - 14x = 11 x - 14x + 26 = 11 -13x + 26 = 11 -13x = 37 x= -2.84 ?

OpenStudy (campbell_st):

remember you are replace -2y so the equation is x + -2y = 11 then x + (-13 + 7x) = 11 so then after distributing you get x + 7x - 13 = 11 or 8x - 13 = 11 now just solve for x